OpenStudy (daltonw):

I need help with a question all i really need to know is how to begin it!! 7. Factor completely 5a2 + b. Prime a(5a + b) b(5a2) ab(5a + 1)

OpenStudy (anonymous):

the answer is prime a(5a + b) = 5a^2 + ab b(5a^2) = 5a^2b b(5a+ 1) = 5a^2b + ab

OpenStudy (daltonw):

wow thanks! you didn't have to answer the whole thing! i really appreciate it though!! thanks so much!!! (: i am giving you a medal

OpenStudy (anonymous):

Thanks! (:

OpenStudy (daltonw):

no thank you!! (: